OpenStudy (anonymous):

What are the solutions of the equation? c^2-10c=0

hartnn (hartnn):

you can take c common to give c*(c-10)=0...now c=0,c-10=0...so c=0,10

OpenStudy (anonymous):

C^2 - 10C = 0 C ( C - 10 ) = 0 either C = 0 or C = 10

OpenStudy (anonymous):

Thanks guys :)